Assume that the daily demand for ice-cream is Q = 500 (P - 2) and Q = 11 000 - 1,000P. If the maximum sale price of ice cream is $ 4, how much will be the, producer and consumer profit ?
1
Expert's answer
2020-01-07T12:38:08-0500

Equilibrium price:

500(P-2) = 11000 – 1000P

500P - 1000 = 11000 – 1000P

1500P = 12000

P = 12000/1500 = $8

Equilibrium quantity:

Q = 500(8-2) = 3,000


Producer profit:

(8-4)(3,000)(1/2) = 6,000

Consumer profit (less):

(4-8)(3,000)(1/2) = -6,000